2. Using the table of constant, solve the friction of a 1,500 kg car if the driver applies the brakes on a concrete road (x), (a) on dry road


(b), on wet road

1
Expert's answer
2022-01-03T09:01:18-0500

(a)

"F_{fr}=\mu_k mg=0.7\times1500\ kg\times9.8\ \dfrac{m}{s^2}=10290\ N."


(b)

"F_{fr}=\mu_k mg=0.3\times1500\ kg\times9.8\ \dfrac{m}{s^2}=4410\ N."